OpenStudy (anonymous):

which two values of x satisfy the equation radical 3-2 cosx=2

OpenStudy (anonymous):

\[\sqrt{3-2\cos(x)}=2\]

OpenStudy (anonymous):

a medal will be given!!!!!!

OpenStudy (anonymous):

\[3\cos 2(x)+2\sin(x)+1=0\]